OpenStudy (anonymous):

A 0.5 kg ball moves in a circle that is 0.4 m in radius at a speed of 4.0 m/s. calculate it's centripetal acceleration @pooja195

OpenStudy (anonymous):

ac=v^2/r is the formula right? because I have a velocity and radius.

pooja195 (pooja195):

ac = v2 / r v = velocity r = radius of circle Distance around = Circumference C = 2πr T = time for one full cycle (period) v = 2πr / T

OpenStudy (anonymous):

I think its 40m/s^2

pooja195 (pooja195):

4.0(4.0)/0.4)=16/0.4=40 correct!!

OpenStudy (anonymous):

then is the centripetal force 20?

pooja195 (pooja195):

yes

OpenStudy (anonymous):

I think I'm getting the hang of this!! :D

pooja195 (pooja195):

YAY!!!!